40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Mar 6 12:34:18 2021, gautam, Update, LSC, Sensing matrix settings messed with sensMat.png
    Reply  Sun Mar 7 19:56:27 2021, Anchal, Update, LSC, Sensing matrix settings messed with 
Message ID: 15874     Entry time: Sat Mar 6 12:34:18 2021     Reply to this: 15876
Author: gautam 
Type: Update 
Category: LSC 
Subject: Sensing matrix settings messed with 

To my dismay, I found today that somebody had changed the oscillator frequencies for the sensing matrix infrastructure we have. The change happened 2 days and 2 hours ago (I write this at ~1230 on Saturday, 3/6), i.e. ~1030am on Thursday. According to the elog, this is when Anchal and Paco were working on the interferometer, but I can find no mention of these settings being changed. Not cool guys 😒 .

This was relatively easy to track down but I don't know what else may have been messed with. I don't understand how anything that was documented in the elog can lead to this weird doubling of the frequencies.

I have now restored the correct settings. The "sensing matrix" I posted last night is obviously useless.

Attachment 1: sensMat.png  96 kB  | Hide | Hide all
sensMat.png
ELOG V3.1.3-